OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−4, 0) and (−5, 0) (−4, 0) and (5, 0) (4, 0) and (−5, 0) (4, 0) and (5, 0)

OpenStudy (phi):

first, notice you can rewrite \[y+20= x^2-x \] as \[y= x^2-x-20 \] the x-intercepts happen when y is zero. So the question is asking "what x's make this 0 ?" \[x^2-x-20 =0 \] can you factor this? (that is how you find what x's make it zero)

OpenStudy (phi):

or, you can try the different possible solutions given, and see which pair give 0 when you replace x with a number in the equation